OpenStudy (anonymous):

The windows of a downtown office building are arranged so that each floor has six fewer windows than the floor below it. If the ground floor has 52 windows, how many windows are on the eighth floor?

OpenStudy (anonymous):

so, let's see, 52 is to the first floor, 46 is to the second floor so it wud be simple, it wud be 52-(7*6) agreed?

OpenStudy (anonymous):

I dont know what you mean. I learned by the formula a^n = a^1 + (n-1)d and i get all the open question just the word problems get me and I got like -50 which are not even close to the possible anwsers

OpenStudy (anonymous):

@mickey4691

OpenStudy (anonymous):

^ those symbols are actually meant for them to lowered.. once again i dont know how to describe it to you.

OpenStudy (anonymous):

are those like arithmentic progression?

OpenStudy (anonymous):

sequence.

OpenStudy (anonymous):

yea, i got it,

OpenStudy (anonymous):

let's say, x is the number of such a term we want to find, then our equation is [x= a - d(n-1)] since we're dealing with less a decrease in the sequence. a is then the first term d is the common difference and n is the number of floor we're dealing wit, better?

OpenStudy (anonymous):

is it like 4 or 6?

OpenStudy (anonymous):

I put the equation/numbers in the formula.. Im doing something wrong

OpenStudy (anonymous):

show me how yu put it

OpenStudy (anonymous):

if you are just supposed to get the number of windows just subtract 6 windows from each floor... 1st.... 52 windows 2nd.... 46 3rd .... 40 4th.... 34 5th ... 28 6th.... 22 7th .... 16 8th.... 10

OpenStudy (anonymous):

@zonazoo , we're dealing with an equation, not a simple subtraction. anyone can do that

OpenStudy (anonymous):

okay... well then just put it into the equation windows = 58 - 6x

OpenStudy (anonymous):

where x is the floor number

OpenStudy (anonymous):

what you were doing is assuming that 52 was the starting point, but its not... it is where x = 1... you want where x = 0, which is at 58

OpenStudy (anonymous):

did that make sense?

OpenStudy (anonymous):

i showed here this let's say, x is the number of such a term we want to find, then our equation is [x= a - d(n-1)] since we're dealing with less a decrease in the sequence. a is then the first term d is the common difference and n is the number of floor we're dealing wit, better? buh she didnt respond

OpenStudy (anonymous):

so 10?

OpenStudy (anonymous):

i think yall are just making this more complicated than it really is... this is just a linear function... and like all linear functions they just need a slope and a y intercept... the slope is fairly easy to see, since it is decreasing by 6 each floor it is -6... and since we want the y intercept or the place where x = 0, that is at 58, so the equation is just y = 58 - 6x... and yes... it is 10

OpenStudy (anonymous):

na, she's dealing with http://www.mathsisfun.com/algebra/sequences-sums-arithmetic.html

OpenStudy (campbell_st):

if you use the foormula, a - 52, d = -6 and n = 8 \[T_{8} = 52 + (8 -1) \times -6\] that will give the answer..

OpenStudy (anonymous):

there is no difference in what I am saying and what you are talking about... if you are doing a linear equation it is just y = 58 - 6x if you are doing arithmetic progression the formula is the same a(n) = a(1) + (n-1)d so ... a(8) = 52 + (8 - 1) - 6 which is a(8) = 52 + 7 (-6) = 10

OpenStudy (anonymous):

both ways it is the same thing... in arithmetic you are starting at the 1st floor, and using a linear equation you are basically starting where x = 0.

OpenStudy (anonymous):

same answer either way... but if you are required to use progression then it would be a(8) = 52 + (8-1)-6

OpenStudy (anonymous):

Thank you Both (:

OpenStudy (anonymous):

u are welcome... anytime.